# Multimodal Inverse Attention Network with Intrinsic Discriminant Feature Exploitation for Fake News Detection

### v1

@@ 1 Introduction
Figure 1: Typical types of fake news. (a)/(b) fabricated fake news where either the text or image is authentic while the other is manipulated to create fake content. Specifically, in (a), Objects in the news image simultaneously feature a fish’s body alongside a pig’s ears and snout, highlighting categorical inconsistencies between regions; The overall semantics of the news text conflict with specific phrases when placed in the realistic context. (c) mismatched text and images from unrelated real news sources, creating inconsistent narratives.
Research on fake news detection has evolved from relying exclusively on text-based approaches Yu et al. (2017); Wu et al. (2024) to incorporating multimodal content, driven by the increasing prevalence of multimedia content on social media platforms Yu et al. (2022). However, the lack of a systematic categorization of news content poses a substantial barrier to advancing effective multimodal fake news detection methods. To address this limitation, this study introduces a novel perspective that classifies multimodal fake news into two distinct categories: a) fabricated fake news encompasses cases where textual and visual content appear consistent but involve deliberate manipulations or synthetic generation in one or more modalities. It presents a more formidable detection challenge, particularly given the rapid advancements in generative AI technologies Zhao et al. (2023); Li et al. (2023), as demonstrated in Figures 1 (a) and (b); b) mismatched fake news refers to instances where textual and visual content are semantically or logically inconsistent. These inconsistencies constitute a form of misinformation capable of distorting public perception and influencing behavior, as illustrated in Figure 1 (c).
@@ 2.1 News Content-based Methods
The lack of correlation between textual and visual content in news is a key characteristic of certain types of multimodal fake news. Consequently, several studies have focused on measuring multimodal consistency to verify the credibility of news. Zhou et al. Zhou et al. (2020) utilized a pre-trained image captioning model to convert images into text which achieved cross-modal semantic space alignment, then assessed the multimodal consistency between the original text and the generated image captions. Similarly, Xue et al. Xue et al. (2021) employed shared weights to enforce the alignment of textual and visual representations within a shared semantic space, calculating the similarity of the transformed multimodal representations. To mitigate the semantic gap between modalities, Jiang et al. Jiang et al. (2023) leveraged CLIP Radford et al. (2021) to extract features from news text and images, subsequently using cosine similarity to guide multimodal fusion. Chen et al. Chen et al. (2022) introduced cross-modal contrastive learning as an auxiliary task, and then used the Kullback-Leibler (KL) divergence to measure the ambiguity score between the latent distributions of text and image sampled from the autoencoder. Qi et al. Qi et al. (2021) measured multimodal entity inconsistency by calculating the similarity between entities in the news text and the corresponding visual content.
In the fusion of multimodal information, the news text and images are not simply modeled through operations such as addition or concatenation. Instead, their high-level semantic interactions are leveraged as crucial features, capturing the intricate relationships between modalities. Qian et al. Qian et al. (2021) extracted text features at different hierarchical levels and fused them with visual features using a contextual transformer to learn cross-modal contextual features and supplementary information. Wu et al. Wei et al. (2022) extracted spatial and frequency domain features from images and progressively fused them with textual features through stacked cross-attention mechanisms Lu et al. (2019). However, their approaches primarily focus on exploring inter-modal relationships, often neglecting the role of inherent modal properties in fake news detection. Moreover, capturing high-level semantic interactions based on latent consistency cues between modalities is typically performed in a black-box manner.
@@ 2.2 External Knowledge-based Methods
Due to limitations in text length or the emergence of novel terms, discrepancies in understanding news content may arise. As a result, several studies have incorporated external knowledge extracted from social media Zhang et al. (2021), knowledge graphs Zhang et al. (2019); Dun et al. (2021) or internet retrieval Zhang et al. (2024a); Qi et al. (2024) to enhance the performance of fake news detection. Zheng et al. Zheng et al. (2022) employed an attention mechanism to combine social graph features, composed of user and comment data, with textual and visual features of news, generating discriminative features for fake news detection. Wu et al. Wu et al. (2023) performed the deep fusion of text and image features based on three distinct reading patterns while leveraging the relationship between news comments and content to capture semantic inconsistencies. Zhang et al. Zhang et al. (2023) leveraged news text entities as prompts to guide a large-scale vision-language model in generating image entities that enhance the semantic knowledge of the content. Zhang et al. Zhang et al. (2024b) converted news images and externally relevant knowledge into pure textual representations, which were then combined with the original text and fed into a prompt-based large language model to predict the authenticity of the news. While some methods enhance the representation of news content by incorporating external knowledge, this simultaneously introduces the risk of noise, which is crucial for the accuracy of fake news detection. These challenges significantly hinder the progress of multimodal fake news detection.
@@ 3 Methodology
3.1 Overview
News content-based fake news detection aims to leverage both textual and associated visual information to evaluate the authenticity of news. Given a dataset of fake news detection $D=\{X,Y\}$, where $X=[{X^{T}},{X^{V}}]$ represents the text ${X^{T}}$ and images ${X^{V}}$ components of the news $X$ and $Y\in\{0,1\}$ denotes the corresponding label indicating whether the news is Fake (0) or Real (1). The objective of this model is to learn a mapping function $F:({X^{T}},{X^{V}})\to Y$ on the training instances of $D$, aiming to map the input space $({X^{T}},{X^{V}})$ to the label vector $Y$. Subsequently, based on the test instances, $F$ is utilized to predict the corresponding labels.
Our MIAN aims to fully learn intrinsic discriminant feature both in intra- and inter-modality for fake news detection, with the detailed architecture illustrated in Figure 2. We first extract unimodal embeddings from the news text and images using modality-specific encoders. For multimodal news, the text content is represented as a text embedding with ${m}$ tokens ${\boldsymbol{T}=\{{t_{1}},{t_{2}},...,{t_{m}}\}}$, ${\boldsymbol{T}\in{\mathbb{R}^{m\times d}}}$, where each token ${{t_{i}}\in{\mathbb{R}^{d}}}$ is a ${d}$-dimensional vector learned from pre-trained BERT Devlin et al. (2018). Meanwhile, we employ pre-trained ViT Dosovitskiy et al. (2020) to learn a visual embedding of image content and gain ${u}$ outputs $\boldsymbol{O}=\left\{{{o_{1}},{o_{2}},...,{o_{u}}}\right\}$, ${\boldsymbol{O}\in{\mathbb{R}^{u\times d}}}$. We then present a hierarchical learning module with two blocks, the Local-to-Global Block and the Local-to-Local Block, which enables multi-granularity interaction to enhance unimodal representation for text and image contents respectively, the detail description is in Section 3.2. Subsequently, MIAN fuses the enhanced multimodal representations using a co-attention mechanism in the cross-modal interaction module, the more detailed description is given in Section 3.3. Additionally, to enable the model to explicitly learn inconsistency features, we design the inverse attention mechanism and integrate it into both the hierarchical learning module and the cross-modal interaction module, the specifics are discussed in Section 3.4. Finally, the enhanced modality-specific representations acquired from the hierarchical learning modules in both the text and image branches, along with the fused representation from the cross-modal interaction module, are concatenated to form the final news representation, which is then fed into the classifier to evaluate the news authenticity.
@@ 3.2 Hierarchical Learning Module (HLM)
To comprehensively explore the internal relationships within news text and image representations and produce enhanced unimodal features, we propose the Hierarchical Learning Module (HLM). After acquiring modality-specific embeddings from pre-trained encoders for text and images, HLM further models intra-modal associations at different levels from two aspects: the Local-to-Local Block, which captures contextual interactions among local features, and the Local-to-Global Block, which establishes dependency relationships between local and global features.
Local-to-Local Block. We employ a multi-head attention mechanism Vaswani et al. (2017) to learn consistent relationships between tokens and capture intrinsic structural features, thereby generating context-aware refined local representations within each unimodal branch. Specifically, for the text modality, given the text embedding $\boldsymbol{T}$, we first preform different multi-head linear transformations to derive the query, key, and value matrices: ${\boldsymbol{Q}_{t}^{i}}=\boldsymbol{TW}_{ll-q}^{i}$, ${\boldsymbol{K}_{t}^{i}}=\boldsymbol{TW}_{ll-k}^{i}$ and ${\boldsymbol{V}_{t}^{i}}=\boldsymbol{TW}_{ll-v}^{i}$, and then concatenate the outputs from the attention mechanism.
$\boldsymbol{Att}_{t}^{i}=softmax\left({\frac{{{\boldsymbol{Q}}_{t}^{i}{\boldsymbol{K}}{{{}_{t}^{i}}^{\top}}}}{{\sqrt{{d_{k}}}}}}\right)$ (1)
$\boldsymbol{h}_{t}^{i}=\boldsymbol{Att}_{t}^{i}{\boldsymbol{V}_{t}^{i}}$ (2)
$\boldsymbol{\bar{R}}_{t}=\left[{h_{t}^{1},...,h_{t}^{n}}\right]{\boldsymbol{W}_{cat}}$ (3)
@@
$\boldsymbol{\hat{R}}_{t}^{ll}=LN(\boldsymbol{T}+\boldsymbol{W}_{t}^{fc1}\boldsymbol{\bar{R}}_{t})$ (4)
$\boldsymbol{R}_{t}^{ll}=LN\left({\boldsymbol{\hat{R}}_{t}^{ll}+ReLU\left({\boldsymbol{W}_{t}^{fc2}\boldsymbol{\hat{R}}_{t}^{ll}+b}\right)}\right)$ (5)
@@
Similarly, the ViT encoder transforms the segmented image into a sequential representation of local features, enabling the application of the self-attention mechanism. Following a similar process as described above, we derive $\boldsymbol{R}_{o}^{ll}$.
Local-to-Global Block. This block refines local features by integrating global context, using the global features to bridge fine-grained local information with overarching semantics. Specifically, for the visual modality, given the image embedding $O$, we construct the global feature, which encapsulates the overall semantics of the image, from two complementary perspectives. On the one hand, ${g^{mean}}={1\mathord{\left/{\vphantom{1n}}\right.\kern-1.2pt}n}\sum\limits_{i=1}^{n}{{o_{i}}}$, which summarizes the spatial information of the entire image. On the other hand, we use the $\left[{cls}\right]$ token ${g^{cls}}$ output from the modality-specific encoder, which summarizes the semantic information of the target objects in the image. Finally, these two representations are concatenated to form the global feature.
${g_{o}}=\left[{{g^{mean}},{g^{cls}}}\right]$ (6)
In this block, the global features act as the query guidance vector, while the local features function as the matched vectors. Their inner product is computed to quantify the similarity between the global and local features. The resulting similarity scores are normalized using a softmax function, ensuring that the attention weights across all regions sum to 1. Specifically, the attention weights for consistency between the global high-level semantic features and the word-level semantic features of the text are computed using the following formula:
$h_{o}=tanh(\boldsymbol{W}_{o}^{1}{\boldsymbol{O}}\odot tanh(\boldsymbol{W}_{o}^{2}{g_{o}}))$ (7)
$att_{o}^{lg}=softmax\left({\boldsymbol{W}_{o}^{3}h_{o}}\right)$ (8)
$\boldsymbol{R}_{o}^{lg}=att_{o}^{lg}{\boldsymbol{O}}$ (9)
where all $\boldsymbol{W}$ are learnable parameters, $h_{o}$ represents the hidden state of the text attention function, and $\odot$ represents the elementwise multiplication. The process of calculating the weighted relationship $\boldsymbol{R}_{t}^{\lg}$ between global and local information in the news text branch is similar to that in the image branch.
Finally, we concatenate the enhanced modality-specific features output by the Local-to-Local Block and Local-to-Global Block in each modality branch to form the hierarchical enhanced representations. Specifically, for the text modality, the representation is ${\boldsymbol{R}_{t}}=\left[{\boldsymbol{R}_{t}^{ll},\boldsymbol{R}_{t}^{\lg}}\right]$, and for the visual modality, it is ${\boldsymbol{R}_{o}}=\left[{\boldsymbol{R}_{o}^{ll},\boldsymbol{R}_{o}^{\lg}}\right]$.
@@ 3.3 Cross-modal Interaction Module (CIM)
To fuse multimodal representations, we propose the Cross-modal Interaction Module (CIM), which leverages co-attention mechanism to capture inter-modal dependencies and effectively integrate the hierarchical enhanced representations of news text and images. Specifically, to obtain the text features enriched with image information, given the hierarchical enhanced representations $\boldsymbol{R}_{t}$ and $\boldsymbol{R}_{o}$ from the HLM, we also apply different multi-head linear transformations $\boldsymbol{Q}_{t}^{co}=\boldsymbol{R}_{t}\boldsymbol{W}_{t}^{co-q}$, $\boldsymbol{K}_{t}^{co}=\boldsymbol{R}_{t}\boldsymbol{W}_{t}^{co-k}$ and $\boldsymbol{V}_{t}^{co}=\boldsymbol{R}_{t}\boldsymbol{W}_{t}^{co-v}$ to derive the query, key, and value matrices, respectively. Due to the similarity of operations with the Local-to-Local Block and space constraints, the specific details of the multi-head attention mechanism are omitted here.
$\boldsymbol{Att}_{t\to o}^{co}=softmax\left({\frac{{\boldsymbol{Q}_{t}^{co}\boldsymbol{K}{{{}_{t}^{co}}^{\top}}}}{{\sqrt{{d_{k}}}}}}\right)$ (10)
$\boldsymbol{\bar{R}}_{t\to o}^{co}=\boldsymbol{Att}_{t\to o}^{co}{\boldsymbol{V}_{t}^{co}}$ (11)
$\boldsymbol{\hat{R}}_{t}^{co}=LN({\boldsymbol{R}_{t}}+\boldsymbol{W}_{co}^{fc1}\boldsymbol{\bar{R}}_{t\to o}^{co})$ (12)
$\boldsymbol{R}_{t}^{co}=LN\left({\boldsymbol{\hat{R}}_{t}^{co}+ReLU\left({\boldsymbol{W}_{co}^{fc2}\boldsymbol{\hat{R}}_{t}^{co}+b}\right)}\right)$ (13)

### v2

@@ 1 Introduction
Figure 1: Typical types of fake news. (a)/(b) fabricated fake news where either the text or image is authentic while the other is manipulated to create fake content. Specifically, in (a), the subject in the image features a fish’s body alongside a pig’s ears and snout, highlighting categorical inconsistencies between regions; In (b), the overall semantics of the news text conflict with specific phrases when placed in the realistic context. (c) mismatched text and images from unrelated real news sources, creating inconsistent narratives.
Research on fake news detection has evolved from relying exclusively on text-based approaches Yu et al. (2017); Wu et al. (2024) to incorporating multimodal content, driven by the increasing prevalence of multimedia content on social media platforms Yu et al. (2022). However, the lack of a systematic categorization of news content poses a substantial barrier to advancing effective multimodal fake news detection methods. To address this limitation, this study introduces a novel perspective that classifies multimodal fake news into two distinct categories: a) fabricated fake news encompasses cases where textual and visual content appear consistent but involve deliberate manipulations or synthetic generation in one or more modalities. It presents a more formidable detection challenge, particularly given the rapid advancements in generative AI technologies Zhao et al. (2023); Li et al. (2023), as demonstrated in Figure 1 (a) and (b); b) mismatched fake news refers to instances where textual and visual content are semantically or logically inconsistent. These inconsistencies constitute a form of misinformation capable of distorting public perception and influencing behavior, as illustrated in Figure 1 (c).
@@ 2.1 News Content-based Methods
The lack of correlation between textual and visual content in news is a key characteristic of certain types of multimodal fake news. Consequently, several studies have focused on measuring multimodal consistency to verify the credibility of news. Zhou et al. Zhou et al. (2020) utilized a pre-trained image captioning model to convert images into textual descriptions which achieved cross-modal semantic space alignment, then assessed the multimodal consistency between the original text and the generated captions. Similarly, Xue et al. Xue et al. (2021) employed shared weights to enforce the alignment of textual and visual representations within a unified semantic space, calculating the similarity of the transformed multimodal representations. To mitigate the semantic gap between modalities, Jiang et al. Jiang et al. (2023) leveraged CLIP Radford et al. (2021) to extract features from news text and images, subsequently using cosine similarity to guide multimodal fusion. Chen et al. Chen et al. (2022) introduced cross-modal contrastive learning as an auxiliary task, and then used the Kullback-Leibler (KL) divergence to measure the ambiguity score between the latent distributions of text and image sampled from the autoencoder. Qi et al. Qi et al. (2021) measured multimodal entity inconsistency by calculating the similarity between entities in the news text and the corresponding visual content.
Under limited available information, some methods refine unimodal features and employ hierarchical cross-attention mechanisms Lu et al. (2019) to model inter-modal relationships. Qian et al. Qian et al. (2021) extracted text features at different hierarchical levels and fused them with visual features using a contextual transformer to learn cross-modal contextual features and supplementary information. Wu et al. Wei et al. (2022) extracted spatial and frequency domain features from images and progressively fused them with textual features through stacked cross-attention mechanisms. However, their approaches primarily focus on exploring inter-modal relationships. Even though a few methods attempt to enhance unimodal feature representations, they overlook intra-modal interactions, which are crucial for detecting fake news, particularly fabricated news. Moreover, most existing methods model high-level semantic interactions through implicit pseudo-consistency, making it challenging to explicitly capture cross-modal inconsistencies.
@@ 2.2 External Knowledge-based Methods
Due to limitations in text length or the emergence of novel terms, discrepancies in understanding news content may arise. As a result, several studies have incorporated external knowledge extracted from social media Zhang et al. (2021), knowledge graphs Zhang et al. (2019); Dun et al. (2021) or internet retrieval Zhang et al. (2024a); Qi et al. (2024); Yu et al. (2020) to enhance the performance of fake news detection. Zheng et al. Zheng et al. (2022) employed an attention mechanism to combine social graph features, composed of user and comment data, with textual and visual features of news, generating discriminative features for fake news detection. Wu et al. Wu et al. (2023) performed the deep fusion of text and image features based on three distinct reading patterns while leveraging the relationship between news comments and content to capture semantic inconsistencies. Zhang et al. Zhang et al. (2023) leveraged news text entities as prompts to guide a large-scale vision-language model in generating image entities that enhance the semantic knowledge of the content. Zhang et al. Zhang et al. (2024b) converted news images and externally relevant knowledge into pure textual representations, which were then combined with the original text and fed into a prompt-based large language model to predict the authenticity of the news. While some methods enhance the representation of news content by incorporating external knowledge, this simultaneously introduces the risk of noise, which is crucial for the accuracy of fake news detection. These challenges significantly hinder the progress of multimodal fake news detection.
@@ 3 Methodology
Fake news detection based on multimodal content aims to leverage both textual and corresponding visual information to evaluate the authenticity of news. Given a dataset of fake news detection $D=\{X,Y\}$, where $X=[{X^{T}},{X^{V}}]$ consists of the textual component $X^{T}$ and the visual component $X^{V}$, and $Y\in\{0,1\}$ is the corresponding label indicating whether the news is fake (0) or real (1). The objective of the model is to learn a mapping function $F:({X^{T}},{X^{V}})\to Y$ from the training set of $D$, which maps the input space $({X^{T}},{X^{V}})$ to the labels $Y$. The learned function $F$ is then used to predict labels for the test instances.
[…]
@@ 3.1 Hierarchical Learning Module (HLM)
To comprehensively explore the internal relationships within news text and image representations and to produce enhanced unimodal features, we propose the Hierarchical Learning Module (HLM). After obtaining modality-specific embeddings from pre-trained encoders for text and images, HLM further models intra-modal associations at different levels from two perspectives: the Local-to-Local Block, which captures contextual interactions among local features, and the Local-to-Global Block, which establishes dependency relationships between local and global features.
Local-to-Local Block. We employ a multi-head attention mechanism Vaswani et al. (2017) to learn consistent relationships between tokens and capture intrinsic structural features, thereby generating context-aware, refined local representations within each unimodal branch. Specifically, for the text modality, given the text embedding $\boldsymbol{T}$, we first perform different multi-head linear transformations to derive the query, key, and value matrices: ${\boldsymbol{Q}_{t}^{i}}=\boldsymbol{TW}_{ll-q}^{i}$, ${\boldsymbol{K}_{t}^{i}}=\boldsymbol{TW}_{ll-k}^{i}$ and ${\boldsymbol{V}_{t}^{i}}=\boldsymbol{TW}_{ll-v}^{i}$, and then concatenate the outputs from the attention mechanism.
$\boldsymbol{Att}_{t}^{i}=softmax\left({\frac{{{\boldsymbol{Q}}_{t}^{i}{\boldsymbol{K}}{{{}_{t}^{i}}^{\top}}}}{{\sqrt{{d_{k}}}}}}\right),$ (1)
$\boldsymbol{h}_{t}^{i}=\boldsymbol{Att}_{t}^{i}{\boldsymbol{V}_{t}^{i}},$ (2)
$\boldsymbol{\bar{R}}_{t}=\left[{h_{t}^{1},...,h_{t}^{n}}\right]{\boldsymbol{W}_{cat}},$ (3)
@@
$\boldsymbol{\hat{R}}_{t}^{ll}=LN(\boldsymbol{T}+\boldsymbol{W}_{t}^{fc1}\boldsymbol{\bar{R}}_{t}),$ (4)
$\boldsymbol{R}_{t}^{ll}=LN\left({\boldsymbol{\hat{R}}_{t}^{ll}+ReLU\left({\boldsymbol{W}_{t}^{fc2}\boldsymbol{\hat{R}}_{t}^{ll}+b}\right)}\right),$ (5)
@@
Similarly, the ViT encoder transforms the segmented image into a sequential representation of local features, enabling the application of the self-attention mechanism. Following the same procedure as described above, we derive $\boldsymbol{R}_{o}^{ll}$.
Local-to-Global Block. This block refines local features by integrating global context, using global features to bridge fine-grained local information with overarching semantics. Specifically, for the visual modality, given the image embedding $O$, we construct the global feature to encapsulate the overall semantics of the image from two complementary perspectives. On the one hand, the mean-pooled vector ${g^{mean}}={1\mathord{\left/{\vphantom{1n}}\right.\kern-1.2pt}n}\sum\limits_{i=1}^{n}{{o_{i}}}$ summarizes the spatial information of the entire image. On the other hand, we use the $\left[{cls}\right]$ token ${g^{cls}}$ output from the modality-specific encoder to represent the semantic information of the salient objects in the image. Finally, these two representations are concatenated to form the global feature:
${g_{o}}=\left[{{g^{mean}},{g^{cls}}}\right].$ (6)
In this block, the global features serve as the query guidance vector, while the local features function as the matched vectors. Their inner product is computed to measure the similarity between the global and local features. The resulting similarity scores are normalized via a softmax function, ensuring that the attention weights across all regions sum to 1. Specifically, the attention weights reflecting the consistency between the global high-level semantic features and the word-level semantic features of the text are computed as follows:
$h_{o}=tanh(\boldsymbol{W}_{o}^{1}{\boldsymbol{O}}\odot tanh(\boldsymbol{W}_{o}^{2}{g_{o}})),$ (7)
$att_{o}^{lg}=softmax\left({\boldsymbol{W}_{o}^{3}h_{o}}\right),$ (8)
$\boldsymbol{R}_{o}^{lg}=att_{o}^{lg}{\boldsymbol{O}},$ (9)
where all $\boldsymbol{W}$ are learnable parameters, $h_{o}$ denotes the hidden state of the text attention function, and $\odot$ represents the element-wise multiplication. The process of calculating the weighted relationship $\boldsymbol{R}_{t}^{\lg}$ between global and local information in the news text branch follows the same procedure as in the image branch.
Finally, we concatenate the enhanced modality-specific features produced by the Local-to-Local Block and Local-to-Global Block in each modality branch to obtain the hierarchically enhanced representations. Specifically, for the text modality, the representation is ${\boldsymbol{R}_{t}}=[{\boldsymbol{R}_{t}^{ll},\boldsymbol{R}_{t}^{\lg}}]$, and for the visual modality, it is ${\boldsymbol{R}_{o}}=[{\boldsymbol{R}_{o}^{ll},\boldsymbol{R}_{o}^{\lg}}]$.
@@ 3.2 Cross-modal Interaction Module (CIM)
To fuse multimodal representations, we propose the Cross-modal Interaction Module (CIM), which leverages a co-attention mechanism to capture inter-modal dependencies and effectively integrate the hierarchically enhanced representations of news text and images. Specifically, to obtain the text features enriched with image information, given the hierarchical enhanced representations $\boldsymbol{R}_{t}$ and $\boldsymbol{R}_{o}$ from the HLM, we apply distinct multi-head linear projections $\boldsymbol{Q}_{t}^{co}=\boldsymbol{R}_{t}\boldsymbol{W}_{t}^{co-q}$, $\boldsymbol{K}_{t}^{co}=\boldsymbol{R}_{t}\boldsymbol{W}_{t}^{co-k}$ and $\boldsymbol{V}_{t}^{co}=\boldsymbol{R}_{t}\boldsymbol{W}_{t}^{co-v}$ to derive the query, key, and value matrices, respectively. Due to the similarity of operations to the Local-to-Local Block and space constraints, the specific details of the multi-head attention mechanism are omitted here.
$\boldsymbol{Att}_{t\to o}^{co}=softmax\left({\frac{{\boldsymbol{Q}_{t}^{co}\boldsymbol{K}{{{}_{t}^{co}}^{\top}}}}{{\sqrt{{d_{k}}}}}}\right),$ (10)
$\boldsymbol{\bar{R}}_{t\to o}^{co}=\boldsymbol{Att}_{t\to o}^{co}{\boldsymbol{V}_{t}^{co}},$ (11)
$\boldsymbol{\hat{R}}_{t}^{co}=LN({\boldsymbol{R}_{t}}+\boldsymbol{W}_{co}^{fc1}\boldsymbol{\bar{R}}_{t\to o}^{co}),$ (12)
$\boldsymbol{R}_{t}^{co}=LN\left({\boldsymbol{\hat{R}}_{t}^{co}+ReLU\left({\boldsymbol{W}_{co}^{fc2}\boldsymbol{\hat{R}}_{t}^{co}+b}\right)}\right).$ (13)